# Hamiltonian simulation-based quantum-selected configuration interaction for large-scale electronic structure calculations with a quantum computer

### v1

@@ IV Results and Discussion
IV.1 The spin-singlet ground (S0) state and the first excited spin-triplet (T1) state of oligoacenes
@@
The results of the HSB-QSCI simulations of oligoacenes are shown in Figure 4. The RHF and ROHF-like single configurational wave functions are used as the starting wave functions for the Hamiltonian simulations of the S0 and T1 states, respectively. The number of Slater determinants in the CAS-CI wave function is 104 (S0) and 61 (T1) for benzene, 15912 (S0) and 11076 (T1) for naphthalene, and 2945056 (S0) and 2255121 (T1) for anthracene, in the point group. Our numerical simulations revealed that the Slater determinants that are important to describe the target electronic states are efficiently sampled from the Hamiltonian simulations. In benzene, the system size is very small and all possible Slater determinants are sampled up to the 7th steps. The is nonzero even if all Slater determinants are included in the QSCI, due to numerical discrepancy, specifically, the different program used for the CAS-CI and the QSCI. We used the eigsh function in SciPy [45] for the CAS-CI energy calculations, while the QSCI is performed using the selected CI module in PySCF [46]. Importantly, chemical precision () was achieved with three steps of the Hamiltonian simulations in all three molecules studied. In anthracene with three steps of the time evolutions, the HSB-QSCI considers 23607 and 22935 determinants for S0 and T1, respectively, which corresponds to 0.80% and 1.02% of the total determinants. The number of Slater determinants considered in the HSB-QSCI calculations is smaller than that in the CISDTQ calculation within the same active space (91601 and 66852 for S0 and T1 states, respectively), but the HSB-QSCI energy is much lower than the CISDTQ values. The values of the CISDTQ with (14e,14o) active space were calculated to be 5.17 and 18.80 kcal mol-1 for the S0 and T1 states, respectively. Interestingly, we found that up to 7-electron excited determinants were sampled in a Hamiltonian simulation with in anthracene. The ability of Hamiltonian simulations to capture higher-order excitations is astounding. These results illustrate the ability of the Hamiltonian simulation to sample a polynomial number of important Slater determinants from exponentially large Hilbert space. The error in the energy can be systematically reduced by increasing the time steps. We have also studied the dependence of the HSB-QSCI energy on the number of shots in anthracene. By increasing the number of shots to , the number of Slater determinants sampled from the Hamiltonian simulation increases by a factor of about 1.5, and the HSB-QSCI energy decreases systematically (see Figure S7 in Supporting Information).
IV.2 Open-shell spin-singlet and excited spin-triplet states of phenylene-1,4-dinitrene
@@
Phenylene-1,4-dinitrene has two major resonance structures, the diradical with a quinonoid skeleton as described in Figure 2, and the dinitrene with an aromatic ring. Its electronic ground state is an open-shell spin-singlet state with two unpaired electrons in in-plane 2p orbitals of the nitrogen atoms. The thermally excited spin-triplet state of phenylene-1,4-dinitrene has been observed by ESR spectroscopy [47, 48, 49], and the singlet–triplet energy gap is determined to be 0.82 kcal mol-1 from the Curie analysis of the ESR signal [50]. The experimentally determined zero-field splitting parameter in the excited spin-triplet state is = 0.171 cm-1 [47], which is about 10 times larger than the value estimated from a point-dipole approximation. This deviation was explained by the lack of electron correlation effect in the point-dipole approximation [51], insisting that sophisticated consideration of electronic correlation is essential to describe its electronic structure.
@@
In the CASSCF optimized orbital basis, the CAS-CI wave function of the electronic ground state is mainly described by the HF configuration () and the HOMO–LUMO two-electron excited configuration from the HF configuration () (see Table S13 in Supporting Information). In the HSB-QSCI simulations of the open-shell spin-singlet ground state, we performed Hamiltonian simulations with three different initial wave functions, , , and the CAS-CI(2e,2o) wave function , to investigate the dependence of the initial wave function on the energies and the convergence behavior of the HSB-QSCI method. The QSCI simulations were performed under four different conditions: (1) use the results with shots for each time step, (2) use the with shots, (3) use the samples from with shots, and (4) merge the samples from and with shots for each step. For the lowest spin-triplet state we used the ROHF-like single determinant as the initial wave function for the Hamiltonian simulation. The HSB-QSCI results are summarized in Figure 5. Our numerical simulations suggest that almost the same number of Slater determinants are sampled when and are used. Interestingly, the number of Slater determinants sampled from the Hamiltonian simulation is smaller when is used as the initial wave function, while the accuracy of the HSB-QSCI energy remains almost unchanged. This result suggests that the probability of sampling less important Slater determinants is reduced by using more accurate initial wave functions. The accuracy of the HSB-QSCI energies of the spin-singlet ground state and the first excited spin-triplet state is comparable, and the singlet–triplet energy gap is calculated to be 1.07 kcal mol-1, in good agreement with the CAS-CI value (1.06 kcal mol-1) and the experimental one [50].
IV.3 The spin-singlet ground state and the first excited spin-triplet state of hexa-1,2,3,4,5-pentaene
@@
Our last example is the spin-singlet ground state and the first excited spin-triplet state of hexa-1,2,3,4,5-pentaene. The one-dimensional sp-hybridized carbon atom chain molecules are called as carbyne. From the DFT calculations, the carbyne with =CH2 termination is predicted to have a helical conjugation in the non-planar geometries [52, 53]. In the lowest spin-singlet state, the planar structure is the energy minimum, but in the lowest spin-triplet state, the planar geometry is a saddle point and the twisted geometry becomes stable. As a result, the singlet–triplet energy gap strongly depends on the dihedral angle between two =CH2 terminations. In this study, we calculated the lowest spin-singlet and the first excited spin-triplet states of the geometries with the dihedral angles of and .
@@
Because carbyne is a one-dimensional molecule, using LMOs and adopting Hamiltonian truncation by maximum locality may be a good option to reduce the computational cost of Hamiltonian simulation. In this work, we investigated Hamiltonian term truncation by operator locality of the Pauli string in the qubit Hamiltonian. Since LMOs are reordered by relative distances before constructing a qubit Hamiltonian, operator locality-based truncation is roughly equivalent to Hamiltonian term truncation based on spatial distances. To assess the effect of Hamiltonian truncation with the locality, we first calculated the number of Hamiltonian terms in the second quantized Hamiltonian and the fidelity of the ground state wave function, , where and are the ground state wave functions with the truncated and untruncated Hamiltonians, respectively. The results are summarized in Tables 1 and 2 for planar and twisted geometries, respectively.
@@
Note that the number of Hamiltonian terms is larger in the twisted geometry than in the planar one, which is due to the difference in the spatial point group. The planar geometry belongs to , which has eight irreducible representations, while the twisted geometry is calculated in point group with four irreducible representations. As can be seen from Tables 1 and 2, the number of Hamiltonian terms is reduced by about 40% when considering up to 10-local terms, with the fidelity of the ground state wave function being about 0.99. The rapid decrease of the fidelity for the maximum locality smaller than 10 can be explained by the fact that the – and – interactions (here we assumed that one-dimensional carbon atom chain is parallel to the z-axis) with the next nearest neighbor C=C bond (for example, – in Figure 3) are described as 10-local operators in the qubit Hamiltonian.
@@
The results of the HSB-QSCI simulations are shown in Figure 6. In the first excited spin-triplet state calculations, we set the starting wave function to carry two unpaired electrons in the central bonds ( and in Figure 3). It is clear that the number of Slater determinants sampled from the Hamiltonian simulations and the convergence behavior of the HSB-QSCI energies are almost the same for , where is the maximum locality of the qubit Hamiltonian terms. These results are consistent with the trend of the singlet ground state fidelity value. The singlet–triplet energy gap calculated using the real-time evolution with the untruncated Hamiltonian () is 53.93 and 20.26 kcal mol-1 for planar and twisted geometries, respectively, and the difference from the CAS-CI singlet–triplet energy gap is less than 0.01 kcal mol-1. When the maximum locality is set to be = 6 or 8, the error in the HSB-QSCI energy becomes larger, but the HSB-QSCI still achieves chemical precision by considering 10 time steps. The calculated singlet–triplet energy gaps are 54.16 (, planar), 54.04 (, planar), 20.38 (, twisted), and 20.26 (, twisted) kcal mol-1. As a result of partial cancellation of the error in the total energies of the singlet and triplet states, the error in the singlet–triplet energy gap is at most 0.23 kcal mol-1.
@@
It should be emphasized that the ROHF-like single configurational wave function in the LMO basis used in the first excited spin-triplet state calculation has rather a small overlap with the CAS-CI wave function. The squared overlap values in the first excited spin-triplet state are 0.1691 and 0.1240 for planar and twisted geometries, respectively (see Tables S16 and S18 in Supporting Information for the CAS-CI wave function). The fact that the HSB-QSCI provides accurate energy with such a small overlap is promising, because the overlap between the HF and the full-CI wave functions decreases with system size, and the preparation of a sophisticated approximate wave function becomes challenging for larger molecules.
IV.4 Hardware demonstration of the spin-singlet ground state and the first excited spin-triplet state calculations of hexa-1,2,3,4,5-pentaene
@@
Since the Hamiltonian term truncation based on the locality of the qubit Hamiltonian terms works excellently in hexa-1,2,3,4,5-pentaene, we performed the HSB-QSCI calculation on the IBM superconducting quantum processor. To reduce the quantum circuit depth of the Hamiltonian simulation, we performed the MPO-based classical compression of the quantum circuit. This approach was demonstrated in our previous study on the quantum phase difference estimation-based energy gap calculations [9]. The details of the MPO-based quantum circuit optimization are given in the reference [9]. In this work, we used a Hamiltonian truncated by the maximum locality ( = 6), and the 6-layer brick wall type quantum circuit is generated for the time evolution operator of . The number of sweeps on the optimizations was set to 10000. In constructing the MPOs, the second-order Trotter decomposition with a single Trotter slice was employed, and the singular value cutoff for the SVD was set to . The order of terms in the Trotter decomposition is set as follows. First, we generate the truncated second quantized Hamiltonian by checking the one- and two-electron integrals in the lexicographic order: after in , and , , , , in that order in . Then the truncated second quantized Hamiltonian is transformed into a qubit Hamiltonian while retaining the order of the terms, and the Trotterized time evolution operator is constructed subsequently. The number of shots was set to be , and the HSB-QSCI was performed five times with and without the SCCR procedure. We also ran the HSB-QSCI calculations with shots to see the shot number dependence.
@@
We first checked the Hamming weight distribution of the bit strings obtained in the measurements and compared it to the uniform distribution. The Hamming weight distributions at are plotted in Figure 7. We confirmed that the Hamming weight distributions of the bit strings obtained from the Hamiltonian simulation with are significantly different from the uniform distribution. However, due to hardware noise, only 19–26% of the measured bit strings have correct Hamming weights (10 in this case).
The HSB-QSCI results are summarized in Figure 8. The differences of the HSB-QSCI energies without SCCR using shots for each time step from the CAS-CI values are about an order of magnitude larger than the statevector results. However, the error in the energy is systematically reduced by using the SCCR. The error of the HSB-QSCI energies is larger for the spin-triplet state than for the spin-singlet ground state, presumably due to the strong multiconfigurational character in the spin-triplet state with the LMO basis. The accuracy of the HSB-QSCI energies improved as the number of shots increased. In fact, when we performed shots for each time step, the values are calculated to be 0.0046, 0.1482, 0.0015, and 0.0406 kcal mol-1 for planar (spin-singlet), planar (spin-triplet), twisted (spin-singlet), and twisted (spin-triplet), respectively, achieving chemical precision.
Since the lowest spin-triplet state of hexa-1,2,3,4,5-pentane shows a strong multiconfigurational character in the LMO basis, we investigated the feedback from the HSB-QSCI results to improve the initial wave function. Using the QSCI wave function calculated from the real-time evolution data up to , we selected three Slater determinants with the CI coefficient and constructed the three-configurational wave functions, by maintaining the ratio of the coefficients. The resulting wave functions are and . Here, we used the orbital ordering in Figures S5 and S6 for planar and twisted geometries, respectively, to represent the wave function. Since the preparation of requires non-nearest neighbor two-qubit gates, the depth of the quantum circuit increased from 22 to 102, in the twisted geometry with . By using the dynamical decoupling [54] in the quantum circuit execution, the Hamming weight distribution after ten steps of time evolution shows a clear difference from the uniform distribution (see Figure S8 in Supporting Information). Interestingly, the HSB-QSCI with SCCR gave very accurate energy in at most 3 steps (see Figure S9 in Supporting Information). These results strongly support the usefulness of the HSB-QSCI for accurate quantum chemical calculations on a quantum computer.

### v2

@@ IV Results and Discussion
Subsections A, B, C are devoted to show numerical simulations for the S0 and T1 states of oligoacenes, phenylene-1,4-dinitrene, and hexa-1,2,3,4,5-pentaene, respectively. Subsection D gives a hardware demonstrations for the S0 and T1 states of carbyne molecules.
IV.1 The S0 and T1 states of oligoacenes
@@
The results of the HSB-QSCI simulations of oligoacenes are shown in Figure 4. The RHF and ROHF-like single configurational wave functions are used as the starting wave functions for the Hamiltonian simulations of the S0 and T1 states, respectively, and the number of shots for each time step was set as . The number of Slater determinants in the CAS-CI wave function is 104 (S0) and 61 (T1) for 1, 15912 (S0) and 11076 (T1) for 2, and 2945056 (S0) and 2255121 (T1) for 3, in the point group. Our numerical simulations revealed that the Slater determinants that are important to describe the target electronic states are efficiently sampled from the Hamiltonian simulations. In 1, the system size is very small, and all possible Slater determinants are sampled up to the fifth steps. In 2, chemical precision () was achieved with three steps of the Hamiltonian simulations. In 3, we found that 10 steps of time evolution with shots are not enough to achieve chemical precision. We expect that the number of important Slater determinants increases with , which is the same as the scaling of Hamiltonian terms, when the target state wave function is well approximated by the single determinant as in the ground state of oligoacenes and the energy gap between the ground and the first excited states is unchanged. Considering the fact that the energy gap is smaller in 3 than in 2, we need more than 3.81 times more shots than in the calculation of 2 to obtain the HSB-QSCI energy of 3 with the same accuracy. We have also studied the dependence of the HSB-QSCI energy on the number of shots and the number of time steps in 3. By increasing the number of shots from and simulating more time steps, the number of Slater determinants sampled from the Hamiltonian simulation increases, and the HSB-QSCI energy decreases systematically (see Figure S14 in Supporting Information). Although we need more steps and shots to achieve chemical precision for larger systems, it should be noted that up to 7-electron excited determinants were sampled in a Hamiltonian simulation with in 3. The ability of Hamiltonian simulations to capture higher-order excitations is astounding. These results illustrate the ability of the Hamiltonian simulation to sample a polynomial number of important Slater determinants from exponentially large Hilbert space. When the kernel_fixed_space subroutine in PySCF is used for subspace Hamiltonian diagonalization, chemical precision was achieved with only three steps of time evolutions with shots in 3. However, the HSB-QSCI wave function considers about 57% and 46% of the Slater determinants, hence the HSB-QSCI wave function is less compact (see Figure S11 in Supporting Information).
IV.2 The S0 and T1 states of phenylene-1,4-dinitrene
@@
4 has two major resonance structures, the diradical with a quinonoid skeleton as shown in Figure 2, and the dinitrene with an aromatic ring. Its electronic ground state is an open-shell spin-singlet state with two unpaired electrons in in-plane 2p orbitals of the nitrogen atoms. The thermally excited spin-triplet state of 4 has been observed by ESR spectroscopy [48, 49, 50], and the singlet–triplet energy gap is determined to be 0.82 kcal mol-1 from the Curie analysis of the ESR signal [51]. The experimentally determined zero-field splitting parameter in the excited spin-triplet state is = 0.171 cm-1 [48], which is about 10 times larger than the value estimated from a point-dipole approximation. This deviation was explained by the lack of electron correlation effect in the point-dipole approximation [52], insisting that sophisticated consideration of electronic correlation is essential to describe its electronic structure.
@@
In the CASSCF optimized orbital basis, the CAS-CI wave function of the S0 state is mainly described by the HF configuration () and the HOMO–LUMO two-electron excited configuration from the HF configuration () (see Table S17 in Supporting Information). In the HSB-QSCI simulations of the S0 state, we performed Hamiltonian simulations with three different initial wave functions, , , and the CAS-CI(2e,2o) wave function , to investigate the dependence of the initial wave function on the energies and the convergence behavior of the HSB-QSCI method. The QSCI simulations were performed under four different conditions: (1) use the results with shots for each time step, (2) use the with shots, (3) use the samples from with shots, and (4) merge the samples from and with shots for each step. For the T1 state, we used the ROHF-like single determinant as the initial wave function for the Hamiltonian simulation. The HSB-QSCI results are summarized in Figure 5. Our numerical simulations suggest that almost the same number of Slater determinants are sampled when and are used. Interestingly, the number of Slater determinants sampled from the Hamiltonian simulation is smaller when is used as the initial wave function. The accuracy of the HSB-QSCI energy with is about an order of magnitude worse than the HF-reference one, but still chemical precision was achieved in 2 steps of time evolution. The singlet–triplet energy gap is calculated to be 1.11 kcal mol-1, in good agreement with the CAS-CI value (1.06 kcal mol-1) and the experimental one [51]. Using PySCF for the subspace Hamiltonian diagonalization slightly increases the number of Slater determinants and yields lower HSB-QSCI energies. The trends of the initial wave function dependence are similar between PyCI- and PySCF-based implementations (see Figure S12 in the Supporting Information).
IV.3 The S0 and T1 states of hexa-1,2,3,4,5-pentaene
@@
Our last example of numerical simulations is the S0 and T1 states of 5. From the DFT calculations, the carbyne with =CH2 termination is predicted to have a helical conjugation in the non-planar geometries [53, 54]. In the S0 state, the planar structure is the energy minimum, but in the T1 state, the planar geometry is a saddle point and the twisted geometry becomes stable. As a result, the singlet–triplet energy gap strongly depends on the dihedral angle between two =CH2 terminations. In this study, we calculated the S0 and T1 states of the geometries with the dihedral angles of and .
@@
Because carbyne is a one-dimensional molecule, using LMOs and adopting Hamiltonian truncation by maximum locality may be a good option to reduce the computational cost of Hamiltonian simulation. In this work, we investigated Hamiltonian term truncation by operator locality of the Pauli string in the qubit Hamiltonian. By reordering the LMOs by relative distances before constructing a qubit Hamiltonian, operator locality-based truncation is roughly equivalent to Hamiltonian term truncation based on spatial distances. To assess the effect of Hamiltonian truncation with the locality, we first calculated the number of Hamiltonian terms in the second quantized Hamiltonian and the fidelity of the ground state wave function, , where and are the ground state wave functions with the truncated and untruncated Hamiltonians, respectively. The results are summarized in Tables 1 and 2 for planar and twisted geometries, respectively.
@@
Note that the number of Hamiltonian terms is larger in the twisted geometry than in the planar one, which is due to the difference in the spatial distribution of orbitals. In planar geometry, the one-electron MO integrals between the nearest neighbor orbitals (e.g., and in Figure 3) are zero, while in twisted geometry, they are non-zero due to helical conjugations. As can be seen from Tables 1 and 2, the number of Hamiltonian terms is reduced by about 40% when considering up to 10-local terms, with the fidelity of the ground state wave function being about 0.99. The rapid decrease of the fidelity for the maximum locality smaller than 10 can be explained by the fact that the – and – interactions (here we assumed that one-dimensional carbon atom chain is parallel to the z-axis) with the next nearest neighbor C=C bond (for example, – in Figure 3) are described as 10-local operators in the qubit Hamiltonian.
@@
The results of the HSB-QSCI simulations are shown in Figure 6. The number of shots for each time step is . In the T1 state calculations, we set the starting wave function to carry two unpaired electrons in the central bonds ( and in Figure 3). It is clear that the number of Slater determinants sampled from the Hamiltonian simulations and the convergence behavior of the HSB-QSCI energies are almost the same for , where is the maximum locality of the qubit Hamiltonian terms. These results are consistent with the trend of the S0 state fidelity value. The singlet–triplet energy gap calculated using the real-time evolution with the untruncated Hamiltonian () is and kcal mol-1 for planar and twisted geometries, respectively, and the difference from the CAS-CI singlet–triplet energy gap is less than 0.1 kcal mol-1 ( = and kcal mol-1 for planar and twisted geometry, respectively). The number of sampled Slater determinants is larger for twisted geometry than for planar geometry, reflecting the spatial distribution of the orbitals, as discussed above. When the maximum locality is set to be = 6 or 8, the error in the HSB-QSCI energy becomes larger. However, our numerical simulations revealed that chemical precision can be achieved even with the maximum locality , by increasing the number of shots to more than and setting the number of maximum time steps to (see Figure S15 in Supporting Information). Note that when using the PySCF kernel_fixed_space subroutine for subspace Hamiltonian diagonalization, chemical precision was achieved with up to five steps of time evolution with shots (see Figure S14 in Supporting Information), but this is a consequence of the quadratic increase of the Slater determinants in PySCF.
@@
It should be emphasized that the ROHF-like single configurational wave function in the LMO basis used in the T1 state calculation has rather a small overlap with the CAS-CI wave function. The squared overlap values in the T1 state are 0.1691 and 0.1240 for planar and twisted geometries, respectively (see Tables S16 and S18 in Supporting Information for the CAS-CI wave function). The fact that the HSB-QSCI provides accurate energy with such a small overlap is promising because the overlap between the HF and the full-CI wave functions decreases with system size, and the preparation of a sophisticated approximate wave function becomes challenging for larger molecules.
IV.4 Hardware demonstration of the HSB-QSCI calculations of the S0 and T1 states of carbyne molecules
@@
Since the Hamiltonian term truncation based on the locality of the qubit Hamiltonian terms works excellently in 5, we performed the HSB-QSCI calculations of 5, 6, and 7 on an IBM Quantum Eagle processor, namely ibm_kawasaki. To reduce the quantum circuit depth of the Hamiltonian simulation, we performed the MPO-based classical compression of the quantum circuit. This approach was demonstrated in our previous study on the quantum phase difference estimation-based energy gap calculations [9]. The details of the MPO-based quantum circuit optimization are given in the reference [9]. In this work, we used a Hamiltonian truncated by the maximum locality ( = 10), and the 10-layer brick wall type quantum circuit is generated for the time evolution operator of . The number of sweeps on the optimizations was set to 10000. In constructing the MPOs, the second-order Trotter decomposition with a single Trotter slice was employed, and the singular value cutoff for the SVD was set to . The order of terms in the Trotter decomposition is set as follows. First, we generate the truncated second quantized Hamiltonian by checking the one- and two-electron integrals in the lexicographic order: after in , and , , , , in that order in . Then the truncated second-quantized Hamiltonian is transformed into a qubit Hamiltonian while retaining the order of the terms, and the Trotterized time evolution operator is constructed subsequently.
@@
We first checked the Hamming weight distribution of the bit strings obtained in the measurements and compared it to the uniform distribution. The Hamming weight distributions at are plotted in Figure 7. We confirmed that the Hamming weight distributions of the bit strings obtained from the Hamiltonian simulation with are significantly different from the uniform distribution. However, due to hardware noise, only 11–23% of the measured bit strings have correct Hamming weights (10, 14, and 18 for 5, 6, and 7, respectively).
The HSB-QSCI results of 5 are summarized in Figure 8. Here, the Hamiltonian simulations were carried out with , and the number of shots was set from to . The depth of the quantum circuit after transpilation with optimization level = 3 in qiskit is 21, regardless of the evolution time steps. The SCCR [19] was adopted to recover configurations with a correct number of electrons, before constructing the subspace Hamiltonian matrix. In the first five to eight steps of time evolution, the number of Slater determinants sampled increases rapidly, but thereafter the number of additionally sampled Slater determinants becomes gradual. When the HSB-QSCI is performed up to ten steps of time evolution and shots for each time step, the energy error is about 2 kcal mol-1. Increasing the number of shots further improves the energy, and the error in the total energy with shots were 0.37, 0.56, 0.33, and 0.43 kcal mol-1 for planar (S0), planar (T1), twisted (S0), and twisted (T1), respectively, achieving chemical precision.
| System | CAS-CI | HSB-QSCI; PyCI | HSB-QSCI; PySCF | |||
| #Dets | /Hartree | #Dets | #Dets | |||
| 5 (Planar, S0) | 63504 | 52209 | 0.37 | 63504 | 0.00 | |
| 5 (Planar, T1) | 44100 | 40092 | 0.56 | 44100 | 0.00 | |
| 5 (Twisted, S0) | 63504 | 54585 | 0.33 | 63504 | 0.00 | |
| 5 (Twisted, T1) | 44100 | 40944 | 0.43 | 43260 | 0.01 | |
| 6 (Planar, S0) | 11778624 | 6147941 | 2.00 | 11778624 | 0.00 | |
| 6 (Planar, T1) | 9018009 | 5753210 | 2.93 | 8612604 | 0.01 | |
| 6 (Twisted, S0) | 11778624 | 5838894 | 2.71 | 11758041 | 0.00 | |
| 6 (Twisted, T1) | 9018009 | 5440991 | 2.84 | 8889232 | 0.00 | |
| 7 (Planar, S0) | 2363904400 | 22569820 | 14.26 | 1141628944 | 0.31 | |
| 7 (Planar, T1) | 1914762564 | 14732855 | 23.36 | 526263465 | 3.70 | |
| 7 (Twisted, S0) | 2363904400 | 25866058 | 13.97 | 1096536996 | 0.30 | |
| 7 (Twisted, T1) | 1914762564 | 16356089 | 25.79 | 554484410 | 2.40 | |
The dimensions of the Hamiltonian matrix and the total energies of the HSB-QSCI and the CAS-CI methods are summarized in Table 3. We used and performed shots for each time step in the HSB-QSCI for 6 and 7. In 6, the HSB-QSCI considered about 50–64% of the entire Slater determinant, and it calculated the total energy with an error of about 2.0–2.9 kcal mol-1. Although chemical precision for the total energy was not achieved, the HSB-QSCI provides the singlet–triplet energy gap with an error of 0.93 and 0.12 kcal mol-1 for planar and twisted geometries, respectively, of 6. The energy difference between the planar and twisted geometries is also calculated with an error of about 0.72 and 0.09 kcal mol-1 for the S0 and T1 states, respectively. These results exemplify the ability of the HSB-QSCI to accurately calculate energy differences between two electronic states and two geometries. We expect that the total energy can be further improved by increasing the number of shots in the Hamiltonian simulation. In 7, the error of the HSB-QSCI energy is about 14 kcal mol-1 for the S0 state, and it increases to 23.36 and 25.79 kcal mol-1 for planar and twisted geometries, respectively, of the T1 state. Again, chemical precision for the total energy was not achieved in 7. In addition, the HSB-QSCI overestimated the singlet–triplet energy gap about 9–12 kcal mol-1. As we pointed out above, the CAS-CI wave function of the T1 state in the localized orbital basis shows an inherent multiconfigurational character, and the single Slater determinant used as the initial wave function of the time evolution has small overlap with the CAS-CI wave function. As a result, it is expected that the number of time steps or the number of shots will have to be increased to account for more Slater determinants in order to accurately describe the electronic structure of the T1 state. Nevertheless, it is worth noting that the HSB-QSCI is able to compute the energy difference between the planar and twisted geometries of the S0 state with chemical precision. Note that when PySCF is used for the subspace Hamiltonian diagonalization, the HSB-QSCI provided the total energy of the S0 state with about 0.3 kcal mol-1 of an error, by considering 46–48% of the Slater determinants, but it failed to compute the total energy of the T1 state with chemical precision.
| System | %#Dets222Percentage of the number of Slater determinants calculated as 100 | 333Percentage of the correlation energy in the active space considered in the HSB-QSCI, calculated as 100 |
| 5 (Planar, S0) | 82.21 | 99.99 |
| 5 (Planar, T1) | 90.91 | 99.98 |
| 5 (Twisted, S0) | 85.96 | 99.99 |
| 5 (Twisted, T1) | 92.84 | 99.99 |
| 6 (Planar, S0) | 52.20 | 99.97 |
| 6 (Planar, T1) | 63.80 | 99.96 |
| 6 (Twisted, S0) | 49.57 | 99.96 |
| 6 (Twisted, T1) | 60.33 | 99.96 |
| 7 (Planar, S0) | 0.95 | 99.86 |
| 7 (Planar, T1) | 0.77 | 99.78 |
| 7 (Twisted, S0) | 1.09 | 99.87 |
| 7 (Twisted, T1) | 0.87 | 99.75 |
Table 4 summarizes the percentages of the number of Slater determinants and the correlation energies covered by the HSB-QSCI method in 5–7. It is worth emphasizing that the HSB-QSCI covers more than 99.75% of the correlation energies in the active space in 7, by considering only ca. 1% of all the Slater determinants.